OpenStudy (catpidigm):

Given a slope of 5 and y-intercept of 10, write an equation in slope-intercept form.

OpenStudy (3mar):

May I help?

OpenStudy (catpidigm):

I believe it has something to do with y = mx + b

OpenStudy (catpidigm):

Sure!

OpenStudy (3mar):

I hope that helps!

OpenStudy (catpidigm):

Still stuck.

OpenStudy (3mar):

Where are you stuck?

OpenStudy (catpidigm):

I'm not really sure how to write the equation.

OpenStudy (3mar):

Do you know that your slope (m) is 5 and y-intercept (b) is 10??

OpenStudy (catpidigm):

ohh I wrote it out as 5x + b = 10... so it's actually 5x + 10 = y?

OpenStudy (catpidigm):

or y = 5x + 10?

OpenStudy (3mar):

Right one!

OpenStudy (3mar):

Congratulations! ♥

OpenStudy (catpidigm):

Yay! Thank you!

OpenStudy (3mar):

Thank you for the medal! I appreciate!

OpenStudy (3mar):

Any more questions?